Jorge and Anita, married taxpayers, earn $90,000 in taxable income and $75,000 in interest from an investment in the City of Heflin bonds. Using the U.S. tax rate schedule for married filing jointly, how much federal tax will they owe? What is their average tax rate? What is their effective tax rate? What is their current marginal tax rate?
Answer #1

Only the taxable income of $90,000 should be considered for tax calculation, since the interest income of a City bond is exempted from tax.

[As per 2019 tax bracket]

Federal tax = 19,400 × 10% + (78,950 – 19,400 =) 59,550 × 12% + (90,000 – 19,400 – 59,550 =) 11,050 × 22%

                        = 1940 + 7146 + 2431

                        = $ 11,517 (Answer)

Average tax rate = (Federal tax / Taxable income) × 100

                            = (11,517 / 90,000) × 100

                            = 1,151,700 / 90,000

                            = 12.80% (Answer)

Effective tax = (Federal tax / Total income) × 100

                        = (11,517 / (90,000 + 75,000)) × 100

                        = 1,151,700 / 165,000

                        = 6.98% (Answer)

Marginal tax rate is the latest tax rate on the last dollar of taxable income; this is 22% (Answer)